OpenStudy (anonymous):

what is 2/3 - 3/4 ? Explain answer plz ?

OpenStudy (dumbcow):

find common denominator 4*3 = 12 2/3 - 3/4 = 8/12 - 9/12 = -1/12

jimthompson5910 (jim_thompson5910):

\[\large \frac{2}{3}-\frac{3}{4}\] \[\large \frac{8}{12}-\frac{9}{12}\] \[\large \frac{8-9}{12}\] \[\large -\frac{1}{12}\] So \[\large \frac{2}{3}-\frac{3}{4}=-\frac{1}{12}\]

OpenStudy (anonymous):

what you need is that \[\frac{a}{b}-\frac{c}{c}=\frac{ad-bc}{bc}\]